# Towards Searching for Photons with Energies beyond the PeV Range from Galactic PeVatrons

### v1

@@ 2 Current Measurements in the PeV Range
Table 1: Parameters of the power-law fits following Eq. (1) for the four sources shown in Fig. 1. The values have been taken from [16, 17].
@@ 3 Extrapolating the Energy Spectra
In Fig. 3, $n_{\gamma}$ is shown as a function of the threshold energy for the four sources, following Eq. (2) with the parameters from Tab. 1. Comparing the extrapolations, one finds that the expected number of photons from 1LHAASO J1908+0615u is largest until about $20\,\mathrm{PeV}$, due to the larger value of $\Phi_{0}$. At higher energies, 1LHAASO J2031+4052u* takes over due to the smaller value of $\Gamma$. For both sources, the number of photons is $n_{\gamma}\simeq 0.3\,\mathrm{km^{-2}\,yr^{-1}}$ above $10\,\mathrm{PeV}$, and $0.02$ ($0.006$) $\mathrm{km^{-2}\,yr^{-1}}$ for 1LHAASO J2031+4052u* (for 1LHAASO J1908+0615u) above $100\,\mathrm{PeV}$. The Crab exhibits the steepest spectrum, which leads to the lowest extrapolated number of UHE photons among the four selected sources. Above $10\,\mathrm{PeV}$, $n_{\gamma}$ is about one order of magnitude smaller for the Crab than for 1LHAASO J1908+0615u and 1LHAASO J2031+4052u* (and two orders of magnitude smaller compared to the latter above $100\,\mathrm{PeV}$). For the Galactic Center, the spectral index is similar to 1LHAASO J1908+0615u, but due to $\Phi_{0}$ (here at a pivot energy of $0.025\,\mathrm{PeV}$ instead of $0.05\,\mathrm{PeV}$ as for the other three sources), $n_{\gamma}$ is about a factor ${\sim}5$ below 1LHAASO J1908+0615u.
@@ 4 Prospects for Observing UHE Photons
In the previous section, we calculated $n_{\gamma}$ in the UHE regime for the four sources. To estimate the potential for an actual observation with air-shower experiments, the parameters of the detector have to be taken into account. For a reasonable estimate, mainly the area covered by the experiment and the energy threshold are relevant. Here, we use the parameters of the Pierre Auger Observatory as a numerical example, as it is the largest air-shower experiment in the world. Naturally, the results can be transferred to any current or future air-shower experiment, for example the Telescope Array experiment [32], the Global Cosmic-Ray Observatory GCOS [33]—a next-generation giant air-shower observatory intended to succeed current-generation experiments like the Pierre Auger Observatory—, or the PEPS (Probing Extreme PeVatron Sources) project [34], a proposed specialized instrument targeted at PeV $\gamma$-sources, utilizing a dense array of layered water Cherenkov detectors [35]. We emphasize that the goal here is not to estimate the prospects of observing, e.g., the Crab at ultra-high energies with the Pierre Auger Observatory, but rather the prospects of measuring photons from a source like the Crab, i.e., one that exhibits a similar energy spectrum, with a detector like the Pierre Auger Observatory. The detector could be located anywhere in the world, just as the source could be located anywhere in the sky. For specific sources and detectors, one would have to take into account the visibility of the source from the location of the detector, but we neglect this here. Instead, as mentioned before, we obtain benchmark scenarios from which giant air-shower experiments can determine the sensitivity they should aim for to either observe UHE photons from PeV $\gamma$-sources or provide significant constraints on their flux.
@@
Also listed in Tab. 2 are the expected numbers of (background) cosmic-ray particles (per point source), calculated from the energy spectrum of cosmic rays (see Sec. 3) for the same combinations of detector area and energy threshold and the same measurement time of ten years. These numbers can be compared to the expected numbers of photons. Assuming similar detection efficiencies for photons and cosmic-ray particles, a background suppression to a level of about $10^{-3}$ to $10^{-4}$ is needed to make a discrimination between photons and cosmic-ray particles feasible. Current photon searches at ultra-high energies (and below) already achieve background suppressions to a level of $10^{-3}$ and better [14, 36], and further improvements are likely with improved analysis methods.

### v2

@@ 2 Current Measurements in the PeV Range
Table 1: Parameters of the power-law fits following Eq. (1) for the four sources shown in Fig. 1. The values have been taken from [16, 17]. For the calculation of the benchmark scenarios in this work, only the central values are used. A brief discussion of the impact of varying the spectral indices within their respective uncertainties can be found in Sec. 4.
@@ 3 Extrapolating the Energy Spectra
In Fig. 3, $n_{\gamma}$ is shown as a function of the threshold energy for the four sources, following Eq. (2) with the parameters from Tab. 1. Comparing the extrapolations, one finds that the expected number of photons from 1LHAASO J1908+0615u is largest until about $20\,\mathrm{PeV}$, due to the larger value of $\Phi_{0}$. At higher energies, 1LHAASO J2031+4052u* takes over due to the smaller value of $\Gamma$. For both sources, the number of photons is $n_{\gamma}\simeq 0.3\,\mathrm{km^{-2}\,yr^{-1}}$ above $10\,\mathrm{PeV}$, and $0.02$ ($0.006$) $\mathrm{km^{-2}\,yr^{-1}}$ for 1LHAASO J2031+4052u* (for 1LHAASO J1908+0615u) above $100\,\mathrm{PeV}$. The Crab exhibits the steepest spectrum, which leads to the lowest extrapolated number of UHE photons among the four selected sources. Above $10\,\mathrm{PeV}$, $n_{\gamma}$ is about one order of magnitude smaller for the Crab than for 1LHAASO J1908+0615u and 1LHAASO J2031+4052u* (and two orders of magnitude smaller compared to the latter above $100\,\mathrm{PeV}$). For the Galactic Center, the spectral index is similar to 1LHAASO J1908+0615u, but due to $\Phi_{0}$ (here at a pivot energy of $0.026\,\mathrm{PeV}$ instead of $0.05\,\mathrm{PeV}$ as for the other three sources), $n_{\gamma}$ is about a factor ${\sim}5$ below 1LHAASO J1908+0615u.
@@ 4 Prospects for Observing UHE Photons
In the previous section, we calculated $n_{\gamma}$ in the UHE regime for the four sources. To estimate the potential for an actual observation with air-shower experiments, the parameters of the detector have to be taken into account. For a reasonable estimate, mainly the area covered by the experiment and the energy threshold are relevant. Here, we use the parameters of the Pierre Auger Observatory as a numerical example, as it is the largest air-shower experiment in the world. Naturally, the results can be transferred to any current or future air-shower experiment, for example the Telescope Array experiment [32], the Global Cosmic-Ray Observatory GCOS [33]—a next-generation giant air-shower observatory intended to succeed current-generation experiments like the Pierre Auger Observatory—, or the PEPS (Probing Extreme PeVatron Sources) project [34], a proposed specialized instrument targeted at PeV $\gamma$-sources, utilizing a dense array of layered water Cherenkov detectors [35]. We emphasize that the goal here is not to estimate the prospects of observing, e.g., the Crab at ultra-high energies with the Pierre Auger Observatory, but rather the prospects of measuring photons from a source like the Crab, i.e., one that exhibits a similar energy spectrum, with a detector like the Pierre Auger Observatory. The detector could be located anywhere in the world, just as the source could be located anywhere in the sky. For specific sources and detectors, one would have to take into account the visibility of the source from the location of the detector, but we neglect this here, assuming that the exemplary source is always within the field of view of the exemplary detector. As mentioned before, we therefore obtain benchmark scenarios from which giant air-shower experiments can determine the sensitivity they should aim for to either observe UHE photons from PeV $\gamma$-sources or provide significant constraints on their flux.
@@
Also listed in Tab. 2 are the expected numbers of (background) cosmic-ray particles (per point source), calculated from the energy spectrum of cosmic rays (see Sec. 3) for the same combinations of detector area and energy threshold and the same measurement time of ten years. These numbers can be compared to the expected numbers of photons. Assuming similar detection efficiencies for photons and cosmic-ray particles, a background suppression to a level of about $10^{-3}$ to $10^{-4}$ is needed to make a discrimination between photons and cosmic-ray particles feasible. Current photon searches at ultra-high energies (and below) already achieve background suppressions to a level of $10^{-4}$ and better [14]. We mention here as an example a search for photons with energies above $10^{19}\,\mathrm{eV}$ published by the Pierre Auger Collaboration, using only data from the surface detector array of water Cherenkov detectors [36]. The analysis was applied to a search sample of about $48{,}000$ air-shower events, yielding $16$ candidate events. Assuming these are all background events would lead to an overall background suppression to the level of 3× $3\,{\times}\,10^{-4}$. Naturally, with more complex detectors, a better background suppression can be reached. For example, using additional information from dedicated muon detectors can improve the background suppression to the level of $10^{-5}$ [37].